OpenStudy (anonymous):

use the properties of logarithms to expand the logarithmic function. 1n[(x^2+1)(x-1)]

OpenStudy (anonymous):

ln(x^2+1) + ln (x-1)

OpenStudy (anonymous):

howdid you get that?

OpenStudy (anonymous):

ln(a*b)=lna+lnb

OpenStudy (anonymous):

log(x*y)=logx+logy log(x-y)=logx-logy

OpenStudy (anonymous):

@amir.sat log(x-y)=logx-logy is wrong. it should be log(x/y)=logx-logy

OpenStudy (anonymous):

sorry ur right imade a mistake